# Towards More Realistic Extraction Attacks: An Adversarial Perspective

### v2

@@ 6.3 Data Deduplication
A commonly recommended solution to extraction attacks and memorization is data deduplication, involving the removal of duplicate data entries within a dataset (Carlini et al., 2023). While costly, data deduplication represents a critical aspect of data curation and has been shown to mitigate extraction risks (Carlini et al., 2023). To understand the role of data deduplication in our discussion, we repeat our experiments using the models from Pythia trained on the deduplicated Pile dataset. The results are collected in Figure 5(b).
@@ 7 Case Studies with Stronger Adversary
We conclude by highlighting the value of our stronger adversary in various case studies.
@@ 7.1 Detecting Pre-Training Data
Extraction attacks are a primary tool in identifying whether certain data was included in a model’s training set. This can be valuable in assessing whether a model is trained on proprietary or sensitive data without permission, evaluating data contamination and leakage in various benchmarks, ensuring regulatory compliance to data governance policies, or even academic research to track the influence of datasets on the model.
While membership inference attacks (MIAs) have been commonly used to detect pre-training data, Maini et al. (2024) argues that MIAs are as good as random guessing when it comes to distinguishing between members and non-members from the same distribution. They show that these attacks learn how to distinguish between concepts, and not actual text, highlighting the importance of using IID data of members and non-members to appropriately perform dataset inference.
We borrow their setup and extend it to the composite setting by increasing the size of the training set for learning correlations. Thus, our composite setting can be alternatively seen as an augmentation technique for the training set. We record the $p$-value of the null hypothesis "the dataset was not used for training" for the Pile dataset in Figure 6(a), under different sizes of the original training data.
We find that the $p$-values for the composite setting with prompt lengths are noticeably lower than those for the baseline, especially at smaller dataset sizes. Thus, our adversary requires less data to achieve the same $p$-value as the baseline. The dataset inference setup by Maini et al. (2024) requires obtaining IID data that the data owners are certain was not used for training by the LLM, which can be difficult to find. Hence, reducing the amount of such data required can be extremely useful, which further emphasizes the value of considering a real-world adversary. Interestingly, we did not find similar strong trends for composite attacks across different model checkpoints. We believe this might be because membership inference information can change drastically across models, and thus combining information from multiple checkpoints does not help in learning better correlations.
@@ 7.2 Copyright Infringement
Copyright issues due to LLMs regurgitating their training data have been heavily studied in recent literature. Karamolegkou et al. (2023) discusses different thresholds for quoting a text ad verbatim that has been considered a violation of fair use, for example, $50$ words is a common threshold used for magazine articles, chapters, etc., while $300$ words is a common threshold used for books. The authors suggest using the longest common subsequence (ROUGE-L score length) as a measure of quantifying text reproduction and potential violations.
Following their reasoning, we record the distribution of ROUGE-L lengths for $2000$ randomly chosen examples in Figure 6(b), both for the strongest baseline as well as the composite settings. We find that a real-world adversary generates more potential copyright violations than the adversary in the literature, which highlights the underestimation of such risks. We note that copyright is a highly complex problem, and simply extracting data from the model might not necessarily constitute a copyright violation. However, our focus is on improving the technical underpinnings that are necessary for a fruitful discussion of copyright issues in LLMs.
@@ 7.3 PIIs Extraction Risk
Another commonly studied risk of memorizing training data is extracting personally identifiable information (PIIs). We use the setup of Li et al. (2024) to create our PII extraction test set from the Pile dataset. We use GLiNER (Zaratiana et al., 2024) to detect $2000$ unique PIIs in the Pile dataset, followed by cutting the sentence right before the PII to create the input prompt. These prompts were fed to the model, and the attack is considered successful if the correct PII is generated anywhere within the first 100 tokens, marking the risk of PII leakage (Li et al., 2024).
We record the extraction risk for the best single setup and composite extraction risks across model checkpoints and model sizes. Since the prompts in this setup are of varying lengths, we do not extend our changing prompt lengths setting to this case study. Similar to Definition 3.1, the composite PII extraction is considered successful if the PII is present in the generation of at least one of the models. The results are collected in the table below, and continuing previous trends, we see a noticeable increase in the extraction rate for an adversary with access to multiple checkpoints.
@@
8 Limitations and Future Work
@@
By highlighting the multi-faceted access available to an adversary in the current LLM landscape, our work reveals a severe underestimation of information leakage risks in the existing literature. We emphasize the importance of explicitly considering the adversarial perspective and the composability of information leakage in extraction attacks.
Our real-world adversary is certainly more powerful but also more expensive than the adversary in the existing literature. Unlike our current setup, where we verify extraction using the ground truth, an adversary would need to justify both the cost of additional model generations and the expense of verifying extracted information. Therefore, future research should explore the cost-benefit trade-offs of multi-faceted access, focusing on when these added expenses may outweigh the benefits of new information extracted, particularly as we show diminishing gains with increased points of access.
As most of our analysis focuses on the risks posed by extraction attacks under the lens of discoverable memorization, future research should also explore how our findings translate to other forms of privacy attacks. Finally, our study addresses the threats posed by powerful real-world adversaries but does not propose specific defence methods. Further exploration is needed to navigate the current LLM ecosystem and mitigate the risks posed by these strong adversaries.
@@ Acknowledgments
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, and Google award. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.

### v3

@@ 6.4 Data Deduplication
A commonly recommended solution to memorization is data deduplication, involving the removal of duplicate data entries within a dataset (Carlini et al., 2023). While costly, data deduplication represents a critical aspect of data curation and has been shown to mitigate extraction risks (Carlini et al., 2023). To understand the role of data deduplication in our discussion, we repeat our experiments using Pythia models trained on the deduplicated Pile dataset. The results are collected in Figure 8(b).
@@
5005006006007007008008009009001,0001{,}00000.20.20.40.4 Train Dataset Sizep $p$-value ($\downarrow$) (a) Dataset Inference p
1001001501502002002502503003000%55%1010%1515% ROUGE-L Length Extraction Rate ($\uparrow$) (b) Copyright Infringement
Best Single Setup (Baseline)Composite Prompt LengthsComposite Model SizesComposite Training Steps
Figure 10: (a) $p$-value for dataset inference (lower is better) across different dataset sizes. The results show significant improvement under the composite setup across different prompt lengths. (b) Extraction rate for different ROUGE-L length thresholds, marking potential copyright violations generated by the LLM. Extraction rates with composite setups are consistently higher than the baseline setup.
@@ 7 Case Studies with Stronger Adversary
We conclude by highlighting the value of our stronger adversary through various case studies.
@@ 7.1 Detecting Pre-Training Data
Extraction attacks identify whether certain data was included in a model’s training set. This can be valuable in assessing whether a model is trained on proprietary or sensitive data without permission, evaluating data contamination and leakage in various benchmarks, ensuring regulatory compliance with data governance policies, or even academic research to track data contamination.
While membership inference attacks (MIAs) have been used to detect pre-training data, Maini et al. (2024) argues that in reality, MIAs are as good as random guessing. They show that these attacks learn how to distinguish between concepts, and not actual text, highlighting the importance of using IID data of members and non-members to appropriately perform dataset inference.
We borrow their setup and extend it to the composite setting by increasing the size of the training set for learning correlations. Thus, our composite setting can be alternatively seen as an augmentation technique. We record the $p$-value of the null hypothesis "the dataset was not used for training" for the Pile dataset in Figure 10(a), under different sizes of the original training data.
We find that $p$-values for the composite setting across prompt lengths are noticeably lower than the baseline, especially at smaller dataset sizes. Thus, our adversary requires less data to get the same $p$-value. The dataset inference setup by Maini et al. (2024) requires obtaining IID data, which can be difficult to find. Hence, reducing the amount of such data required can be extremely useful. Interestingly, we did not find strong composition trends across model checkpoints, possibly because membership information can change drastically across models, and thus, combining information from multiple checkpoints might not be helpful.
@@ 7.2 Copyright Infringement
Copyright issues due to LLMs regurgitating their training data have been heavily studied in recent literature. Karamolegkou et al. (2023) discusses different thresholds for quoting a text ad verbatim that can be considered a violation of fair use, for example, a $50$ word threshold for magazine articles, chapters, etc., while a $300$ word threshold for books. The authors suggest using ROUGE-L lengths (longest common subsequence) as a measure of text reproduction and potential violations.
We plot the distribution of ROUGE-L lengths for 2,000 randomly chosen examples in Figure 10(b), comparing the strongest baseline and the composite settings. We find that our adversary produces more potential copyright violations than the baseline, highlighting an underestimation of such risks in existing literature. While copyright law is complex and extraction alone may not imply a violation, our focus is on strengthening the technical underpinnings of copyright issues in LLMs.
@@ 7.3 PIIs Extraction Risk
Another commonly studied risk of memorization is extracting personally identifiable information (PIIs). We use the setup of Li et al. (2024) to create our PII extraction test set from the Pile dataset. We use GLiNER (Zaratiana et al., 2024) to detect $2000$ unique PIIs in the Pile dataset, followed by cutting the sentence right before the PII to create the input prompt. These prompts were fed to the model, and the attack is considered successful if the correct PII is generated anywhere within the first 100 tokens, marking the risk of PII leakage (Li et al., 2024).
We record the extraction risk for the best single setup and composite extraction risks across model checkpoints and model sizes. Since the prompts in this setup are of varying lengths, we do not extend our changing prompt lengths setting to this case study. Similar to Definition 3.1, the composite PII extraction is considered successful if the PII is present in the generation of at least one of the models. The results are collected in Table 1, and continuing previous trends, we see a noticeable increase in the extraction rate for an adversary with access to multiple checkpoints.
@@
Table 1: Composite attacks for PII extraction.
7.4 Closed-Source Models
Our study till now has focused on open-source models due to the availability of training data for verifying the success of our attacks. However, recent work by Duarte et al. (2024) has shown how similar studies can be performed even when a model’s training data is undisclosed.
Duarte et al. (2024) propose DE-COP, a method for detecting memorization in LLMs using a multiple-choice task inspired by counterfactual memorization. The approach presents the target model with four options: one original text passage and three paraphrases generated using a different LLM, in this case, Claude 2. Models tend to choose the original text more frequently if seen during training, thus signaling memorization.
The authors validate this method with a novel benchmark - BookTection, which consists of excerpts from 165 books published both before and after 2023. Since closed-source models released in 2022 could not have been trained on books published after 2023, this serves as a reliable non-member set. The books pre-2023 form the member-set. The method compares a model’s performance on each book to a baseline performance (computed from non-member books), using it to assess whether a specific book was part of the model’s training data. Each book can have multiple passages, so the authors combine the accuracy per passage for the detection of the book.
For our study, we focus on the experiments on LLaMA-2 70B using the BookTection dataset with varying passage lengths. The authors found that shorter passages yield higher F1 scores: 64, 128, and 256-token passages achieve F1 scores of 0.67, 0.65, and 0.64, respectively. Building on this, we extend their method to a composite setting that combines different lengths. This composite significantly boosts detection performance on LLaMA-2 70B, achieving an F1 score of 0.78, underscoring the generalizability of our approach.
8 Discussion
@@
By highlighting the multi-faceted access adversaries have in the current LLM landscape, our work shows that existing literature greatly underestimates information leakage risks, thus emphasizing the importance of explicitly considering the adversarial perspective and the composability of information leakage in extraction attacks. Our work provides a foundation for future exploration and defense against more realistic extraction attacks, contributing to a secure and robust management of the risks associated with memorization in LLMs.
Potential Defenses. We studied the threats posed by real-world adversaries and showed that existing defence methods (such as data deduplication), while undoubtedly useful, are prone to the same risks of composability. However, we did not propose defense techniques to deal with this adversary.
Firstly, it’s important to recognize that not all instances of discoverable memorization are harmful, and therefore may not require a defense. For instance, defending against improved methods of detecting copyright violations or data contamination is inadvisable. Such efforts could hinder those seeking to determine whether their data was misused by companies or developers.
However, when we do want to defend against these attacks in more harmful scenarios, future research could include ways to disrupt multi-faceted access to the dataset. For example, shuffling and re-chunking training sequences for each model can break the link between specific sequences and model behaviour. Since many existing LLM attacks operate at the sequence level (Meeus et al., 2024), this simple randomization can significantly increase the difficulty of combining information across checkpoints or model sizes.
Other defense strategies, such as anonymizing sensitive data before training (Yu et al., 2024) or applying differential privacy (DP) during training (Yu et al., 2022), can also help. However, their effectiveness may decline against a stronger adversary. Finally, beyond modifying training, one may wish to defend already trained models against such attacks. While this is challenging, several practical strategies could be potentially helpful. Output perturbation techniques, such as adding noise or rephrasing responses, can reduce information leakage, even though a determined adversary may still bypass them. Access control measures, such as rate limiting and monitoring for suspicious prompts, also offer practical defenses in real-world deployments.
Beyond Discoverable Memorization. Our analysis focuses on the risks posed by extraction attacks under the lens of discoverable memorization. However, our arguments on the increased privacy surface apply to any adversary with multi-faceted access to the underlying data. Moreover, our experiments on Pythia and OLMo represent a controlled setup where the underlying models were trained on the exact same data and data order. However, in reality, multiple models from the same family might have some differences in their training. Hence, translating our findings to other forms of privacy attacks and dataset homogenization in the real world is an important direction for future research.
@@ Acknowledgments
We thank the anonymous reviewers and the action editor Kai-Wei Chang from TACL, for their continued feedback and comments that helped improve our work. We also thank Pratyush Maini for providing us early access to the dataset inference code (Maini et al., 2024) before its public release.
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Google award, MITACS, FRQNT, and NSERC Discovery Grants program. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.
Appendix A Unique Extraction Results
We provide additional results on the trends of churn across model checkpoints (Figure 11) and prompt lengths (Figure 12). We find similar trends as in the main paper.
1234567890%0.20.2%0.40.4%0.60.6%0.80.8% Number of Checkpoints Extraction Rate (a) Recurrence of Extractions across Training Steps
[…]
100k110k120k130k140k0%0.50.5%11%1.51.5% Training Steps (c) Drop-One Extractions across Training Steps
Figure 11: Granular trends of churn across training steps for Pythia. (a) Number of model checkpoints extracting the same information. Even though the majority of samples are extractable by all 9 checkpoints, a significant amount of extractions are unique to a subset, or even a single model. (b) While later model checkpoints contribute more unique extractions, each model, regardless of size, provides significant unique extractions. (c) Composite extraction rates after dropping a single model checkpoint (x-axis: dropped checkpoint, y-axis: resulting extraction rate) show that extraction rates remain stable even if we remove the contributions of the last checkpoint.
123456789100%0.50.5%11% Number of Prompts Extraction Rate (a) Recurrence of Extractions across Prompt Lengths
[…]
502003505000%11%22%33% Prompt Lengths (c) Drop-One Extractions across Prompt Lengths
Figure 12: Granular trends of churn across prompt lengths for Pythia. (a) Number of prompts extracting the same information. Even though the majority of samples are extractable by all 10 prompt lengths, a significant amount of extractions are unique to a subset, or even a single prompt length. (b) While longer prompt lengths contribute more unique extractions, each prompt length provides some unique extractions. (c) Composite extraction rates after dropping a single prompt length (x-axis: dropped prompt length, y-axis: resulting extraction rate) show that extraction rates remain stable even if we remove the contributions of the longest prompt length.